Four resistance each of R ohms and two resistances each of S ohms are connected (as shown in Fig.



2.50) to four terminasl AB and CD. A p.d. of V votls is applied across the terminals AB and a



resistance

1
Expert's answer
2021-11-22T19:33:31-0500

To be given in question

Resistance

"R_1=R_2=R\Omega"

"R_3=R_4=S\Omega"

"R_1 andR_2" Parallel combination

"R'=\frac{R_1 R_2}{R_1+R_2}"

"R'=\frac{R\times R}{R+R}=\frac{R}{2}\Omega"

Now

"R_3 and R_4 parallel combination"

"R''=\frac{R_3R_4}{R_3+R_4}"

"R''=\frac{S \times S}{S+S}=\frac{S}{2}"


"R_{eq}=R'+R''"

"R_{eq}=\frac{R}{2}+\frac{S}{2}"

"R_{eq}=\frac{R+S}{2}"